Question 8 Find the sum of the following four vectors ((a) and (b) for x and y components respectively), and as (c) a magnitude and (d) an angle relative to +x (including sign plus or minus) P: 22.0 m, at 25.0° counterclockwise from +x : 15.0 m, at 13.0° counterclockwise from ty R: 25.0 m, at 28.00 clockwise from y S: 5.50 m, at 44.0° counterclockwise from -y (a) Number (b) Number (c) Number (d) Number Units Units Units Units
Answer #1

Answers are -

(a) 8.6 m

(b) -2.2m

(c) Magnitude = 8.9 m

(d) Angle = -14.3 deg.เร้ 22.0 kIrite uu ta vectors i, tent vector n-tah…:→ (1937 + g.33), = 凤 -325.oxdn(rao . 2)을 h(Do-2g)3(m + 2.5.o- = (3-81-4.03) -2.2~ -I (-22.) -14-30 = 8.6 Nalus ą thin CV7t rulchve to +x.ax İS =-14.3